OpenStudy (anonymous):

find the value of θ in the equation cos2θ+sin²θ=cos³θ+3θ-6

OpenStudy (anonymous):

i got as far as substituting the double angle formula in there and simplifying to get cos^2x=cos^3x+3x-6 cos^2x-cos^3x=3x-6

OpenStudy (anonymous):

yeah that is how far i got too

OpenStudy (anonymous):

you could try solving \(u^3-u^2+3u-6=0\) and and then take the inverse cosine if you can find such a solution

OpenStudy (anonymous):

okay well its a multiple choice: so x could equal 6 cos2x x or 2

OpenStudy (anonymous):

2 it is then lol thanks!